OpenStudy (anonymous):

Anyone? The height of a triangle is 9 m less than its base. The area of the triangle is 56 m^2. Find the length of the base.

rishavraj (rishavraj):

let length of base be = x so height = x - 9 and area = \[\frac{ 1 }{ 2 }\times base \times height \]

OpenStudy (anonymous):

@rishavraj So you'd only have x-9 * x * 1/2?

rishavraj (rishavraj):

\[\frac{ 1 }{ 2 } \times (x) \times (x - 9) = 56\] solve it and get "x"

OpenStudy (anonymous):

I got -7, and 16

rishavraj (rishavraj):

its 16 and 7 not -7 ....:))))))